OpenStudy (anonymous):

find limit as x approaches -infinity, and infinity y= x^(3/2) (4-9x)

OpenStudy (anonymous):

solve these first\[\lim_{x \rightarrow \infty} x^{3/2}= \] \[\lim_{x \rightarrow \infty} 4-9x=\] \[\lim_{x \rightarrow -\infty} x^{3/2} =\] \[\lim_{x \rightarrow -\infty} 4-9x=\]

OpenStudy (anonymous):

oh you have to seperate them? for positive infinity you get two different answers: infinity and -infiinity. what do you do then?

OpenStudy (anonymous):

(infinity)*(- infinity) =?

OpenStudy (anonymous):

oh gotcha! thanks

OpenStudy (anonymous):

you dont have to seperate them, but it makes life a lot easier

OpenStudy (anonymous):

can i ask you another limit i'm strugglng on